OpenStudy (anonymous):

Calculus help Find the relative extrema of the function f(x) = x^4 - 32x^2 + 3 I've worked this out on the calculator, but I am unsure how to do on pen and paper. I tried completing the square, but it doesn't seem to come out right. Relative maximum at (0, 3); relative minima at (4, -253), (-4, -253)

Vocaloid (vocaloid):

take the derivative

OpenStudy (anonymous):

3x^3-64x This doesn't factor evenly

OpenStudy (anonymous):

\[f \prime \left( x \right)=4x^3-64x\] \[f \prime \left( x \right)=0~gives~4x^3-64x=0\] 4x(x^2-16)=0 x=0,-4,4

OpenStudy (anonymous):

find \[f \prime \prime \left( x \right) ~all~three~ points~for~maxima~and~minima.\]

OpenStudy (anonymous):

Wow, I missed that one, Thanks

OpenStudy (jhannybean):

nope this helps too.

OpenStudy (jhannybean):